OpenStudy (anonymous):

3ro the 2 power + x5

OpenStudy (teddyiswatshecallsme):

What exactly is the question?

OpenStudy (egbeach):

\[3^{2}+5x\] 9+5x=0 5x=-9 x= -9/5 x=1.8 Im assuming this is what the question is but please correct me if i am mistaken.